Grade 11Physical Chemistry

Find out the longest wavelength of light just able to eject a least bound electron from copper. Given work function=4.3 eV.

Krati
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
hc/k = 4.3 + K.E
take K.E. = 0
(for longest wavelength)
therfore,
k = hc/4.3 = 1240/4.3 = 288.3872 nm